Please help me understand how to do this. I need to add these. 1/5 + 3/10 + 3/8 =? The answer comes to 7/8 but not understanding how they got 8 as the common denominator
Answer by rfer(16322)

Get common denominator then reduce.
1/5+3/10+3/8=
8/40+12/40+15/40=35/40=7/8